OpenStudy (lexi724):

MEDAL TO CHECK: MY CHOICE: B The area of circle B is 36 times greater than the area of circle A. If the radius of circle B is 12 feet, what is the radius of circle A? A.2 feet B.4 feet C.6 feet D.36 feet

OpenStudy (igreen):

Find the area of circle B.

OpenStudy (igreen):

\(\sf A = \pi r^2\) \(\sf A = (3.14) (12)^2\) Simplify

OpenStudy (lexi724):

i meant c btw

OpenStudy (igreen):

C is wrong.

OpenStudy (igreen):

Simplify that above there.

OpenStudy (lexi724):

i am but i put down the wrong answer orginally wanted to clear it up anyways 452.16

OpenStudy (igreen):

Yep, now divide it by 36 to get the area for circle A.

OpenStudy (igreen):

'Cause circle B's area is 36 times the area of circle A.

OpenStudy (lexi724):

12.56

OpenStudy (igreen):

Yep, now plug it into the formula to find the area of a circle. \(\sf A = \pi r^2\) Plug in what we know: \(\sf 12.56 = (3.14) r^2\) Divide 3.14 to both sides, what's 12.56 / 3.14?

OpenStudy (lexi724):

4

OpenStudy (lexi724):

2

OpenStudy (igreen):

Square root of 4*

OpenStudy (igreen):

Lol..yes, that's our answer.

OpenStudy (igreen):

\(\sf 4 = r^2\) \(\sf r = \sqrt{4}\) \(\sf r = 2\)